### v1

# Jacobian-Enhanced Neural Networks

Abstract
Jacobian-Enhanced Neural Networks (JENN) are densely connected multi-layer perceptrons, whose training process is modified to predict partial derivatives accurately. Their main benefit is better accuracy with fewer training points compared to standard neural networks. These attributes are particularly desirable in the field of computer-aided design, where there is often the need to replace computationally expensive, physics-based models with fast running approximations, known as surrogate models or meta-models. Since a surrogate emulates the original model accurately in near-real time, it yields a speed benefit that can be used to carry out orders of magnitude more function calls quickly. However, in the special case of gradient-enhanced methods, there is the additional value proposition that partial derivatives are accurate, which is a critical property for one important use-case: surrogate-based optimization. This work derives the complete theory and exemplifies its superiority over standard neural nets for surrogate-based optimization.
Index Terms:
Jacobian, neural network, gradient-enhanced, multi-layer perceptron, surrogate-based optimization, surrogate model, meta-model.Nomenclature
| number of layers in neural network | |
| activation output, | |
| bias (parameter to be learned) | |
| activation function | |
| number of training examples | |
| number of nodes in a layer | |
| number of nodes in input layer, | |
| number of nodes in output layer, | |
| weight (parameter to be learned) | |
| training data input | |
| training data output | |
| predicted output | |
| linear activation input | |
| Hyperparameters | |
| learning rate | |
| hyperparameter to priorities training data values | |
| hyperparameter to priorities training data partials | |
| hyperparameter used for regularization | |
| Mathematical Symbols and Operators | |
| cost function | |
| loss function | |
| first derivative, | |
| second derivative, | |
| Superscripts/Subscripts | |
| node in layer, | |
| input, , | |
| output, , | |
| layer of neural network, | |
| node of current layer, | |
| node of previous layer, | |
| training example, |
I Introduction
Jacobian-Enhanced Neural Networks (JENN) are fully connected multi-layer perceptrons, whose training process is modified to predict partial derivatives accurately. This is accomplished by minimizing a modified version of the Least Squares Estimator (LSE) that accounts for Jacobian prediction error, where a Jacobian extends the concept of a gradient for vector functions with more than one output. The main benefit of jacobian-enhancement (synonymous to gradient-enhancement) is better accuracy with fewer training points compared to standard fully connected neural nets. This paper develops the theory behind JENN and quantifies its benefits for the intended use-case. It complements previous work in the following ways. First, it provides a concise but complete derivation, including equations in vectorized form (see appendix), without relying on automatic differentiation. Second, it provides a deep-learning formulation that can handle any number of inputs, hidden layers, or outputs. Finally, through judicious use of hyperparameters, it enables weighting; i.e. a process by which individual sample points can be prioritized differently in situations that call for it. The latter, in particular, gives rise to the concept of polishing (described in the application section) which offers an additional tuning mechanism to further improve SBO in gradient-descent applications.
I-A Intended Use-Case
JENN is intended for the field of computer aided design, where there is often a need to replace computationally expensive, physics-based models with fast running approximations, commonly referred to as surrogate models or meta-models. Since a surrogate emulates the original model accurately in near-real time, it yields a speed benefit that can be used to carry out orders of magnitude more function calls quickly. This opens the door to useful design applications, such as Monte Carlo simulations of computationally expensive functions for example. However, this is true for any type of surrogate model in general. In the special case of gradient-enhanced methods, there is the additional value proposition that partial derivatives are accurate, which is a critical property for one important use-case: Surrogate-Based Optimization (SBO). The field of aerospace engineering is rich in design applications of such a use-case, with entire book chapters dedicated to it [1, 2] and applied examples commonly found in aerodynamic applications [3, 4, 5].
I-B Limitations
First and foremost, gradient-enhanced methods require factors and responses to be smooth and continuous. Second, the value proposition of a surrogate is that the computational expense of generating training data to fit the model is much less than the computational expense of directly performing the analysis with the original model itself, which is not always true, especially when computing partials. Gradient-enhanced methods are therefore only beneficial if the cost of obtaining partials is not excessive in the first place (e.g. adjoint methods [6] in computational fluid dynamics), or if the need for accuracy outweighs the cost of computing the partials. One should therefore carefully weigh the benefit of gradient-enhanced methods relative to the needs of their application.
I-C Supporting Software
An open-source Python library that implements the theory can be found at https://pypi.org/project/jenn/. There exist many excellent deep learning frameworks, such as tensorflow [7], which are more performant and versatile than JENN, as well as many all-purpose machine learning libraries, such as scikit-learn [8]. However, gradient-enhancement is not inherently part of those popular frameworks, let alone the availability of prepared functions to predict partials after training is complete, and requires non-trivial effort to be implemented in those tools. The present library seeks to close that gap; it is intended for those engineers in a rush with a need to predict partials accurately and who are seeking a minimal-effort Application Programming Interface (API) with low-barrier to entry111Note that GENN from the Surrogate Modeling Toolbox [9] is JENN with a different API (present author contribution).
II Previous Work
A cursory literature research reveals that the field of gradient-enhanced regression is predominantly focused on Gaussian processes. Dedicated book chapters can be found about gradient-enhanced Kriging models [2], salient Python libraries have been almost exclusively devoted to gradient-enhanced Bayesian methods [9], and even review papers about gradient-enhanced meta-models do not mention neural networks, except in passing [10]. However, early work in the late 1990’s can be found [11] which uses gradient-enhanced neural network approximations in the context of Concurrent Subspace Optimization (CSSO), a type of optimization architecture in the field of multi-disciplinary design optimization. Those authors conclude that gradient-enhanced neural networks are more accurate and require significantly less training data than their non-enhanced counterparts, which seems to be the general consensus in all work surveyed for this paper.
Follow-on research by that group provides the mathematical derivation of the equations for neural net gradient-enhancement [12], but the formulation is not general; it is limited to shallow neural networks with only two layers, which is of limited utility. This shortcoming is overcome in the late 2000’s by a deep, neural-network formulation with gradient-enhancement [3], in which the authors apply their algorithm to surrogate-based optimization of airfoil designs using evolutionary algorithms. All necessary equations to reproduce their algorithm are provided, but their formulation is limited to only one output and the vectorized form is missing for computer programming.
Recent work in 2019 takes a different approach, relying instead on established deep-learning frameworks to delegate back-propagation to third-party, automatic-differentiation tools [4]. Those authors successfully apply gradient-enhanced neural networks to aerodynamic use-cases and re-affirm the previous conclusion that such techniques improve accuracy and reduce the amount of training data needed, provided the cost of obtaining the gradient is not counter-productive in the first place. While deep learning frameworks have obvious advantages (e.g. parallel computing), there is complementary research value in developing the equations for back-propagation, which is a special case of automatic differentiation for scalar functions. Stand-alone equations are reproducible without third-party tools and convey deeper insight into the mechanics behind an algorithm, which others can build upon.
Finally, the most recent work on gradient-enhanced neural nets focuses on gradient-enhanced Physics-Informed Neural Networks [13] (PINN). This creative research uses the residuals (and its partials) of Partial Differential Equations (PDE), governing some physical system of interest, as the loss function for training neural nets (instead of using the classic least squares estimator). However, PINN is outside the target scope of the present work, which seeks to develop a general purpose algorithm to fit any smooth function, regardless of what it represents, even if it is not physics-informed.
III Mathematical Derivation
Jacobian-Enhanced Neural Networks (JENN) follow the same steps as standard Neural Networks (NN), shown in Fig. 1, except that the cost function used during parameter update is modified to account for partial derivatives, with the consequence that back propagation must be modified accordingly. This section provides the mathematical derivation for Jacobian-enhancement.
III-A Notation
For clarity, let us adopt a similar notation to the one used by Andrew Ng [14] where superscript denotes the layer number, is the training example, and subscript indicates the node number in a layer. For instance, the activation of the node in layer evaluated at example would be denoted:
To avoid clutter in the derivation to follow, it will be understood that subscripts and refer to quantities associated with the current and previous layer, respectively, evaluated at example without the need for superscripts. Concretely:
Derivatives with respect to input are denoted by the prime symbol and subscript . For example:
In turn, vector quantities will be defined in bold font. For instance, the input and output layers associated with example would be given by:
III-B Normalize Training Data
Normalizing the training data can greatly improve performance when either inputs or outputs differ by orders of magnitude. A simple way to improve numerical conditioning is to subtract the mean and divide by the standard deviation of the data. The normalized data then becomes:
This approach works well, except when all values are very close such that the denominator vanishes as standard deviation goes to zero. In that case, alternative normalization methods (or none) should be considered.
III-C Forward Propagation
Predictions are obtained by propagating information forward throughout the neural network in a recursive manner, starting with the input layer and ending with the output layer. For each layer (excluding input layer), activations are computed as:
where are neural network parameters to be learned. For the hidden layers (), the activation function is taken to be the hyperbolic tangent222Other choices are possible, provided the function is smooth (i.e. not ReLu):
For the output layer (), the activation function is linear:
III-D Parameter Update
The neural network parameters are learned by solving the following, unconstrained optimization problem:
where it should be understood that is short hand for all parameters of the neural network. The cost function is given by the following expression, where is a hyper-parameter to be tuned which controls regularization:
Parameters are updated iteratively using gradient-descent:
where the hyper-parameter is the learning rate. In practice, some improved version of gradient-descent would be used instead, such ADAM [15], but the former is easier to explain without loss of generality. The loss function is taken to be a modified version of the Least Squares Estimator (LSE), augmented with an extra term to account for partial derivatives:
Note that two more hyper-parameters have been introduced: and . They act as weights that can be used to prioritize subsets of the training data in situations that call for it. By default, , meaning that all data is prioritized equally, which is the common use-case. However, in some situations, there may be an incentive to prioritize certain examples over others, as will be shown in the results section. This formulation offers the flexibility to do so.
III-E Backward Propagation
Forward propagation seeks to compute the gradient of the predicted outputs with respect to (abbreviated w.r.t.) the inputs . By contrast, back-propagation computes the gradient of the cost function with respect to the neural net parameters :
III-E1 Loss Function Derivatives w.r.t. Neural Net Parameters
In a first step, assume the quantities and are known and consider the gradient of the loss function with respect to the parameters from layer . Applying the chain rule to Eq. III-D yields:
where the superscript has been dropped for notational convenience, with the understanding that all equations to follow apply to example . The various partial derivatives in Eq. 11 follow from Eq. 3:
Substituting these expressions into Eq. 11 yields:
III-E2 Loss Function Derivatives w.r.t. to Neural Net Activations
All quantities in Eq. 15 are known, except for and which must be obtained recursively by propagating information backwards from one layer to the next, starting with the output layer . From Eq. III-D, it follows that:
The derivatives for the next layer are obtained through the chain rule. According to Eq. III-D, the loss function depends on activations and associated partials , which are themselves functions of and . Hence, according to the chain rule, current layer derivatives w.r.t. to previous layer activations are:
Further applying the chain rule to the intermediate quantities and using Eq. 3, one finds that:
All necessary equations to compute the derivatives of the loss function w.r.t. parameters in any layer are now known. Derivatives are obtained by working our way back recursively across layers, repeatedly applying Eq. 17 after passing the solution in the current layer to the preceding layer as follows:
III-E3 Cost Function Derivatives w.r.t. Parameters
IV Verification and Validation
The theory developed in the previous section was implemented into an installable Python library for which a link was provided earlier. This section shows validation results333code to reproduce available at https://codeocean.com/capsule/9360536/tree for selected test functions in Fig. 2, which compares JENN against a standard Neural Net (NN) when training data is sparse (shown as dots). Higher dimensional examples are available on the project’s homepage, but were not included since it is hard to visualize more than two dimensions on a static plot. All else being equal, it can clearly be seen that JENN outperforms NN in all cases, re-affirming the general concensus that gradient-enhancement achieves better accuracy with less data (provided the cost of obtaining the Jacobian is not counter-productive). Finally, it was also verified that runtime scales as using the Rastrigin function to generate progressively larger datasets and training a new model, as shown in Fig. 2c using a last generation MacBook Pro (2021).
V Example Application: Surrogate-Based Optimization
Surrogate-based optimization is the process of using a fast-running approximation, such as a neural net, to optimize some slow-running function of interest, rather than directly using the function itself. This use-case typically arises when the response of interest takes too long to evaluate relative to the time scale of the application. For example, in real-time optimal control, one may wish to consider a neural network to make quick predictions of complicated physics when updating the trajectory[16]. The purpose of this section is to quantify the benefit of JENN relative to regular neural networks for gradient-based optimization. This will be accomplished by solving the following optimization problem:
| Minimize: | (21) | |||
| With Respect To: | (22) | |||
| Subject To: | (23) | |||
where is a surrogate-model approximation to the famous Rosenbrock function:
Results444code to reproduce available at https://codeocean.com/capsule/9360536/tree are shown in Fig.3 where optimizer convergence histories are shown for the true response and its surrogate-model approximations using JENN and NN under the same settings, except for jacobian-enhancement, respectively. Training data (not shown to avoid clutter) was generate by sampling the true response on a regular grid. Training histories are provided in Figs. 3a–3b to verify that both neural networks converged to the same degree.
V-A Results Without Polishing
In a first step, consider Fig.3c which does not use polishing (explained in the next section). Under perfect conditions, one would expect the optimization history to follow the same path for all models, but this is not observed. For both surrogates, the terminal value is indeed close to zero (the true optimum) but neither JENN nor NN reach the true location of this optimum situated at ; although JENN does get much closer. The Rosenbrock function is notorious for exhibiting a shallow valley with near-zero slope, which creates a significant challenge for surrogate-based optimization. Carefully studying the contour lines, one can indeed observe subtle undulations in this shallow valley within the noise of surrogate prediction error, which create artificial local minima that trap the optimizer.
This finding is not surprising. Recall from Eq. III-D that neural nets learn by minimizing prediction error. Hence, by construction, there is more reward during training for accurately predicting regions of large change, where prediction error implies a large penalty, than there is for accurately predicting small features in the vicinity of the optimum, where prediction error implies negigible penalty. This situation is at odds with the needs of gradient-based optimization, which requires extreme accuracy near the optimum; accuracy elsewhere is desirable but less critical, provided slopes correctly point in the direction of improvement.
V-B Results With Polishing
For most practical engineering problems, close is good enough. However, there are situations where optimization accuracy is desirable if not critical, such as airfoil design (to ensure smooth contours free of undulations). This situation can be ameliorated by using what will be called polishing. It consists of restarting a previously trained model and continue training it after magnifying some portion of the design space to be improved. This is accomplished by judiciously setting or to prioritize specific training points belonging to that subspace. For example, a radial basis function centered on zero could be used to magnify small numbers in the aforementioned shallow valley of the Rosenbrock function:
where and are hyperparameters that control the radius of influence of a point. They were set to and in the results shown. Far away from any point where , the weights are , whereas nears those points . This has the desired effect of disproportionately emphasizing small values of and during training by up to a factor of 1000. The results are shown in Fig. 3d where it can be seen that JENN now recovers the true optimum almost exactly; artificial local minima nearly vanish. This is not true for NN; although results improve noticeably. Overall, this section demonstrates that jacobian-enhancement offer a quantifiable benefit over standard neural nets for SBO applications.
VI Noisy Partials
We now consider the effect of noisy partials. As mentioned in the introduction, many practical design applications rely on legacy software suites, such as missile DATCOM [17] during early design phases, for which analytical or automatically differentiated partials are not available. In this case, the only recourse for computing derivatives is finite difference. However, despite their attractive simplicity, it has been shown that finite difference methods are not as accurate as their counterparts [18]. It is therefore natural to wonder: how accurate must partials be in order to take advantage of JENN in design applications where finite difference is unavoidable?
The answer to this question is shown by example in Fig. 4 using the Rastrigin function as a test bed555code to reproduce available at https://codeocean.com/capsule/9360536/tree once again, where corresponds to baseline results for a standard neural net (gradient-enhancement off). Training data partial derivatives were obtained using finite difference with progressively increasing steps sizes. The corresponding mean finite difference error shown was computed against true analytical partials. It can clearly be seen that beyond seven percent error, the benefit of gradient-enhancement vanishes. In fact, beyond that threshold, partials are too inaccurate to provide useful information: accounting for erroneous partials is worse than not accounting for partials at all. These results are welcome, as they demonstrate that the usefulness of gradient-enhanced methods is not restricted to perfectly accurate partials; even noisy partials can be useful (within a reasonable threshold).
Finally, these results further show that there is no benefit to setting to values other than zero or one; it does not mitigate the effect of noisy partials. For most applications, the hyperparameter should therefore be treated as a binary scalar to turn gradient-enhancement on () or off (); except in special situations which call for distribution (rather than a scalar) such as polishing, as previously illustrated, or incomplete partials. The latter is likely to be a common scenario where partials are not available for all inputs, in which case can be used to turn gradient-enhancement on/off depending on whether a partial is missing or not (i.e. resulting in a tensor of zeros and ones).
VII Conclusion
In summary, this research developed a jacobian-enhanced, deep-learning formulation for dense multi-layer perceptrons and successfully validated the implementation against selected canonical test functions. It was shown that validation results support the general consensus that gradient-enhanced methods require less training data and achieve better accuracy than their non-gradient-enhanced counterparts, provided the cost of obtaining training data partials is not prohibitive in the first place. It was further shown that JENN outperforms standard neural nets in the context of surrogate-based optimization, getting much closer to the true optimum but not quite reaching it for the selected example. In fact, it was observed that a learning procedure which minimizes any version of the least squares estimator is at odds with the needs of gradient descent in the vicinity of the optimum, where slopes are close to zero. However, it was shown that this shortcoming can be overcome through a process termed polishing, which prioritizes salient sample points in that neighborhood through judicious use of hyperparameters. This unique feature enabled by the present formulation provides an additional knob for tuning optimized solution accuracy in situations that call for it. The open-source Python library associated with this work was provided as a link in earlier sections, in the hope that it will be useful to others and spark new contributions in this researcher area.
Acknowledgement
The supporting code for this work used the exercises by Prof. Andrew Ng [14] as a starting point. It then built upon it to include additional features but, most of all, it fundamentally changed the formulation to include jacobian-enhancement and made sure all arrays were updated in place (data is never copied). The author would therefore like to thank Andrew Ng for offering the fundamentals of deep learning on Coursera, which took a complicated subject and explained it in simple terms that even an aerospace engineer could understand.
VIII Key Equations in Vectorized Form
Key equations from the preceding material will now be re-written in a vectorized form (i.e. processing all examples simultaneously without loops) suitable for programming, as a helpful companion to source code. In the notation to follow, the symbol implies element-wise array multiplication.
VIII-A Data Structures
Using capital letters, notation can be extended to matrix notation to account for all training examples, starting with training data inputs and outputs:
The associated Jacobians are formatted as:
Note that the hyperparameters can either be scalars or real arrays of the same shape as and , respectively. The shape of the neural network parameters to be learned (i.e. weights and biases) is independent of the number of training examples. They depend only on layer sizes and are denoted as:
However, hidden layer activations do depend on the number of examples and are stored as:
The associated hidden layer derivatives w.r.t. inputs are stored as:
It follows that hidden layer derivatives w.r.t. a specific input are formatted as:
Finally, the back propagation derivatives w.r.t. parameters are:
VIII-B Forward Propagation (Eq. 3)
VIII-C Parameter Update (Eq. 8)
VIII-D Backward Propagation (Eqs. 15–16, 17, 20)
The loss function partials w.r.t. last layer activations are:
The loss function partials w.r.t. hidden layer activations are:
The loss function partials w.r.t. neural net parameters are:
Finally, the cost function partials w.r.t. neural net parameters are:
where S is used to denote the vectorized summation along the last axis of an array666https://numpy.org/doc/stable/reference/generated/numpy.sum.html. Any other summation shown above implies that it could not be vectorized and required a loop. This only happened when summing over partials. Hence, the number of loops needed equals the number of inputs , which is usually negligible compared to the size of the training data . The latter is the critical dimension to be vectorized.